OpenStudy (anonymous):

find the slope -3x+2y=4

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge y=mx+b$$

OpenStudy (anonymous):

2y=3x+4 y=3/2x + 4/2 y=3/2x + 2 slope is 3/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

always change equation into standard form, m = slope, b = y-intercept